OpenStudy (12wishs):

Can someone help? Zen wants to fill a window box with soil. The length of the window box is 5 feet, width is 6 inches, and height is 9 inches. Each bag of soil contains 0.625 cubic foot of soil. How many bags of soil will Zen need to fill the window box completely? [1 foot = 12 inches]

OpenStudy (12wishs):

I'm thinking its 432

OpenStudy (mathmale):

Make use of the given length, width and depth of the window box to find its VOLUME in cubic inches. Then convert that result to cubic feet. What next? I'd need to know how you obtained 432 (through seeing your work) before responding to y our "I thnk it's 432" post.

OpenStudy (12wishs):

ok i got to it by multiplying 5*6*9=270÷0.625=432

OpenStudy (skullpatrol):

Why are you multiplying feet times inches?

OpenStudy (12wishs):

ohhhh wait that's right my mistake so it would be 60*9*6=3240÷0.625=5184

OpenStudy (skullpatrol):

But 0.625 has the units of cubic FEET?

OpenStudy (12wishs):

yes

OpenStudy (skullpatrol):

Your 3240 has the units of cubic INCHES?

OpenStudy (12wishs):

yes

OpenStudy (12wishs):

yes

OpenStudy (mathmale):

... and so you'll need to convert from cubic inches to cubic feet.

OpenStudy (12wishs):

well I get 332 cubic feet

OpenStudy (12wishs):

432*

OpenStudy (12wishs):

9 bags?

OpenStudy (skullpatrol):

Sounds reasonable :-)

OpenStudy (12wishs):

are you sure?

OpenStudy (skullpatrol):

As a student, it is part of learning to make yourself sure :)

OpenStudy (skullpatrol):

That is the best way of gaining confidence in your calculations.

OpenStudy (12wishs):

I'm pretty sure because that's the lowest it can go without it being a decimal

OpenStudy (12wishs):

no wait since each bag only contains 0.625 cubic feet of dirt it would be 270

OpenStudy (skullpatrol):

What volume did you calculate?

OpenStudy (12wishs):

432 × 0.625=270

OpenStudy (skullpatrol):

6 inches = 0.5 feet 9 inches = 0.75 feet

OpenStudy (mathmale):

Let's do a quick summary of this discussion: The length of the box is 5 feet, or 60 inches. The width is 6 inches. The height (depth) is 9 inches. Volume of the box is (length)(width)(height). Multiply: (60 inches)(6 inches)(9 inches) to obtain the volume in cubic inches. Next, look up or figure out how many cubic inches there are in 1 cubic foot. As before, volume = length * width * height. 1 cubic ft = ? cubic inches

OpenStudy (12wishs):

12 cubic inches

OpenStudy (mathmale):

You can, of course, convert all of the box dimentions into feet, such as 0.5 ft., 0.75 ft and 60 ft. The length of ONE SIDE of a 1-cubic foot cube is 12 inches. What is the AREA of the base of this cube? What is the HEIGHT of this cube? What is the volume of this cube, in cubic feet? Hint: 1 cubic foot is the equivalent of between 1,500 and 2,000 cubic inches. Try again on that one. The number of cubic inches in a cube of side length 12 inches is .... ?? cubic inches.

OpenStudy (mathmale):

We want the volume of the planter in cubic feet, so that we can leave the 0.625 cubic foot volume of the bags of soil as is.

OpenStudy (skullpatrol):

1728 cu in. = 1 cu ft

OpenStudy (mathmale):

Thanks. You said earlier: "12wishs 12 cubic inches". Unfortunately this is not correct. Look at what skullpatrol has posted, above. He got 1,728 cubic inches by multiplying 12 inches times 12 inches times 12 inches. @12wishes? What's your next step? Goal: figure out how many bags of soil are needed to fill this one planter.

OpenStudy (skullpatrol):

Working in feet gives $$\huge\dfrac{5*0.5*0.75}{0.625}$$

OpenStudy (12wishs):

3

OpenStudy (mathmale):

The numerator of this fraction gives us the VOLUME of the planter in cubic feet. The denomiantor of this fraction represents the number of cubic feet in each bag. @skullpatrol?

OpenStudy (mathmale):

@12wishes: 3 what?

OpenStudy (12wishs):

3 bags of dirt

OpenStudy (skullpatrol):

Correct.

OpenStudy (12wishs):

thank you both so much I wish I could give you both a medal

OpenStudy (skullpatrol):

Thanks for asking :-)